# Infinitely many hyperelliptic curves of small genus and small fixed rank, and of any genus and rank two

### v3

@@ 1. Introduction
Let $K$ be a number field, and $C/K$ a nice curve111smooth, projective, geometrically irreducible with its Jacobian $\mathrm{Jac}(C)$. Mordell-Weil theorem states that $\mathrm{Jac}(C)(K)$ is a finitely generated abelian group, i.e., isomorphic to $\mathbb{Z}^{r}\times G$, where $G$ is a finite group. The number $r$ is called the rank of $\mathrm{Jac}(C)$ over $K$.
@@
Even though there are techniques to compute the rank $r$ in certain cases, for example, using descent [Sch95, Sto01, SS03], it is still a difficult problem to determine the exact value of $r$. In principle, the problem can be split into two tasks. The lower bound on $r$ can be obtained from finding enough $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. The upper bound on $r$ can be computed from using descent and computing Selmer groups of $\mathrm{Jac}(C)(K)$. When the two bounds are equal, then the exact value of $r$ can be determined.
@@
  • For studying the set $C(K)$, it is very useful to know the rank of $\mathrm{Jac}(C)(K)$, as in some cases a powerful method of Chabauty and Coleman [Cha41, Col85] can be applied to determine $C(K)$. Some generalisations of the method by Chabauty and Coleman, dependent on the rank of $\mathrm{Jac}(C)(K)$, can be found in [Kim05, Kim09, BDM+19, Sik09]. These methods have led to numerous results on studying modular curves that parametrise certain families of elliptic curves, their modularity, torsion subgroups, isogenies and residual Galois representations.
@@
  • One part of the proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, amounts to prove that there is an abelian variety $A/K$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(A(L))=\mathrm{Rank}(A(K))>0$. This was first proven very recently in the groundbreaking work by Peter Koymans and Carlo Pagano [KP24] who constructed an elliptic curve $E/\mathcal{O}_{K}$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(E(L))=\mathrm{Rank}(E(K))=1$. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [Zyw25b]. A different proof of Hilbert 10th problem using positive rank of Jacobians of some twist families of hyperelliptic curves was shown by another groundbreaking work by Levent Alpöge, Manjul Bhargava, Wei Ho, and Ari Shnidman [ABHS25].
@@
  • The Birch-Swinnerton-Dyer conjecture [BSD65] proposes intricate relations between algebraic properties and analytic properties of abelian varieties. Finding evidence supporting the conjecture has harbored fruitful results in understanding arithmetic properties of elliptic curves [GZ86, Gro91, ZI04, DLR15].
In the case of modular curves, using Eichler-Shimura relations, it is possible to decompose their Jacobians and compute their rank by computing the rank of their components. Still, this technique is not possible for most of the curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. Knowing the existence of additional $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$ may not be helpful, because the 2-Selmer group of $\mathrm{Jac}(C)$ can be of large dimension, which is not sufficient to compute the rank of $\mathrm{Jac}(C)(K)$.
@@
For example, in [Gaj22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large rank. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)-1$ is given in [Gaj22, Theorem 12(1)], whereas $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)$ is given in [Gaj22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [Col85], or Stoll’s improvement [Sto06]. However, giving an exact rank of such curves seems extremely difficult.
In this manuscript, we shift our focus to studying arithmetic statistics of rank of $\mathrm{Jac}(C)(K)$, where $C$ varies over an infinite family of curves. There is a wealth of previous studies which analyse the statistics of rank over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, Smi22a, Smi22b, Smi25], cubic twist families of elliptic curves [ABS22, KS24], quartic twist families of elliptic curves over $\mathbb{Q}(i)$ [Sav], and infinitely many elliptic curves of fixed rank between 0 and 4 over any number field [Zyw25a].
1.1. Our results and strategy
The goal of this manuscript is to construct infinitely many non-isomorphic higher genus curves of fixed positive rank.
Notation. Throughout this manuscript, we will not write explicitly the adjective non-isomorphic and the construction of infinitely many curves will always assume they are non-isomorphic. The proof that they are non-isomorphic in each case is straightforward, because we construct the curves by gluing twists of hyperelliptic curves, and for infinitely many twists, we introduce a new prime of bad reduction (using that we twist by square-free elements).
Motivation and difficulties. To the best of our knowledge, the highest exact rank of an elliptic curve over $\mathbb{Q}$ up to this date is 20, found by Elkies and Klagsburn [EK20] (whereas the same people found and elliptic curve of rank at least 29, which is the highest rank elliptic curve over $\mathbb{Q}$, but it is not known if 29 is the exact rank of this curve). Thus, in principle, we can construct an abelian variety over $\mathbb{Q}$ of dimension $g$ and rank $r=20g$, just take $E^{g}$, where $E$ is the known rank 20 curve.
However, there are great mathematical merits to restrict our attention to Jacobians of curves rather than just random abelian varieties because, as we explained before, the curves play a crucial role in investigating many of the most important mathematical questions, and to study and understand their arithmetic, the knowledge of the rank of its Jacobian is one of the most essential information one can obtain. In particular, even though we admit that it is more difficult to compute the rank of a simple Jacobian, as we also mentioned, studying the modular curves lead to so many exciting and groundbreaking results, and they, in principle, have decomposable Jacobians. Hence, it also motivates studying decomposable Jacobians.
In this manuscript, we pay particular focus to studying Jacobians of hyperelliptic curves. Recall that for $g\geq 2$, the dimensions of the moduli space of principally polarised abelian varieties of dimension $g$, the moduli space of curves of genus $g$, and the moduli space of hyperelliptic curves of genus $g$ over $\mathbb{C}$ are respectively given by $\frac{g(g+1)}{2}$, 3g33g-3, and 2g12g-1. For $g\geq 3$, it is clear that the hyperelliptic locus is a thin subset of both moduli spaces of principally polarised abelian varieties and moduli spaces of genus $g$ curves, so the aforementioned strategy of simply taking products of lower dimensional abelian varieties does not construct Jacobians of hyperelliptic curves. Even for $g=2$, it is not clear whether a product of some two elliptic curves is isogenous to the Jacobian of a hyperelliptic curve over a given number field $K$. These considerations highlight one of the subtleties that this manuscript aims to address: we provide explicit gluing procedure for two hyperelliptic curves to produce a new hyperelliptic curve of higher genusdefined over the given number field $K$, thus focusing our problem to thin subset of moduli space of principally polarised abelian varieties.
Another subtlety lies in ensuring that the gluing procedure allows us to incorporate many deep and technical results regarding constructions of hyperelliptic curves whose Jacobians have small fixed rank. Our article demonstrates that all relevant conditions appearing in such previous works can be checked and satisfied even after undertaking the gluing procedure for two hyperelliptic curves over $K$. In particular, we find examples of curves such that effectively, when gluing them, we can consider one of them as a fixed curve (in fact, we use curves which are isomorphic to that one), and then we show that we can apply the known results about the twists of the other one, or we allow twisting both curves with a good control of the rank of both twists concurrently.
Organisation. Observant readers may realise that Section 2 and Section 3 contain similar arguments regarding gluing two hyperelliptic curves to construct a new hyperelliptic curve whose Jacobian has fixed positive rank. The reason why we introduce the genus two case first before the higher genus case in this manuscript is to present our method to potential readers in a more intuitive and accessible manner. Our contributions to this project initially started with genus two hyperelliptic curves. The gluing method is simpler and easier to state for genus 2 hyperelliptic curves, and we hope introducing these techniques first will allow readers to feasibly follow through the generalisations to cases for higher genus hyperelliptic curves.
In Section 2, we consider bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases. In §2.1, we consider the case over number fields $K$ where both $E_{1}$ and $E_{2}$ do not have any $K$-rational 22-torsion points. In §2.2, we examine elliptic curves over $\mathbb{Q}$ having marked rational $2$-torsion points. In §2.3, we use elliptic curves over $\mathbb{Q}$ having all rational $2$-torsion points. We then construct families of genus two curves $C$ by fixing one curve $E_{2}$ of the prescribed rank $r$, and quadratically twisting $E_{1}$ except in §2.5, which we explain later. This allows us to utilise results from previous literature (such as [Nag29, Fre84, KMR14]) that there are infinitely many such twists of rank 0 or 1 depending on the choice of $E_{1}$. We can hence construct infinitely many bielliptic curves of prescribed rank. For $\boldsymbol{K=\mathbb{Q}}$, we construct the desired families for $\boldsymbol{0\leq r\leq 11}$.
For most of our constructions, it is crucial to fix an elliptic curve of the prescribed rank and to aim for twists of rank 0. Other strategies, for instance, to fix an elliptic curve of rank 0 and to aim for infinitely many twists of rank $r$, seem much more difficult, as in principle, it is much easier to show the existence of infinite families of curves of rank 0 rather than positive rank. We prove that there are not many rational points on these curves, as explained in §2.4. We use positive rank twists in one of the examples in §2.2 and §2.5.
In §2.5, we also construct infinitely many curves over any number field $K$ of genus two, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. These constructions use groundbreaking results in [Smi22b, KP25], which give an existence of infinitely many quadratic twists of a certain elliptic curve over $K$ of rank 0 [Smi22b] or 1 [KP25]. The conditions for elliptic curves for which their theorems can be applied are more complicated than those over $\mathbb{Q}$. Unlike the previous three subsections, the non-zero rank examples we construct crucially depend on rank 1 quadratic twist families. Checking whether the conditions appearing in [KP25] are applicable to our construction of genus two curves is an additional technicality we address in this subsection. The heart of our result for number fields lies in Theorem 2.8, where we propose an intricate construction for such genus two curves. Our work can be considered as a continuation of the recent progress by pioneering constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, Zyw25b]. More concretely, our work is an application of the results from [KP25] and of the same question for rank 0 [Smi22b].
Generalizing our strategy explored for genus two curves, in Section 3 we consider hyperellitpic curves $C$ of genus $g\geq 3$ such that $\mathrm{Jac}(C)$ is isogenous to product of two Jacobians of hyperelliptic curves $C_{1}$ and $C_{2}$. In §3.1, we construct infinitely many curves over any number field $K$ of any genus at least $3$, whose rank of the Jacobian over $K$ is each of 0, 1, and 2. The overall strategy of the proof remains identical, but we employ different previous results to generate such curves. We still use [Yu19, BM, Smi22b] for the existence of infinitely many quadratic twists of Jacobians of a certain hyperelliptic curve over $K$ of rank 0. But we use the recent breakthrough by Peter Koymans and Adam Morgan [KM] for the existence of infinitely many such quadratic twists of rank $1$. We check in §3.1 that the conditions appearing in [Smi22b] and [KM] are once again applicable to our construction of higher genus curves, as we highlight in Theorem 3.10 and Theorem 3.13. In remaining subsections, we give examples of infinitely many hyperelliptic curves over $\mathbb{Q}$ of genus 3, 4, 5, and 6 with fixed positive rank. To the best of our knowledge, our work is the first attempt to utilize gluing method in the setting of generating infinitely many higher genus hyperelliptic curves with positive fixed rank.
The groundbreaking work by Peter Koymans and Adam Morgan [KM] in fact answers the previously posed challenge problem, which asks for generating infinitely many absolutely simple Jacobians of hyperelliptic curves over $\mathbb{Q}$ with rank $1$. We mention the formulation of the previously posed question in Section 4.
1.2. Explicit families of positive rank
Even though there are several instances of families of curves of fixed positive rank, such as [ABS22, KP25, Zyw25b], not so many give examples of families with explicit parametrisation. There are nevertheless some previous studies which give explicit parametrisation of families of elliptic curves over $\mathbb{Q}$ with rank $1$, such as those proven by Monsky [Mon92] and Dasgupta and Voight [DV18]. Also, very recently, the question of whether there are infinitely many rank 2 elliptic curves over $\mathbb{Q}$ was proved by Zywina [Zyw25c], who, impressively, constructs elliptic curves non-isomorphic over $\overline{\mathbb{Q}}$, and we can say that these results are semi-explicit. While the elliptic curves are given by concrete equations, these equations depend on a certain primality condition that can be viewed as non-explicit, but we still believe that this result deserves to be mentioned in this section. We give explicit examples of curves of genus two and a fixed rank between 0 and 4 in §2.2 and §2.3.

### v4

@@ 1. Introduction
Let $K$ be a number field, and $C/K$ a nice curve111smooth, projective, geometrically irreducible of genus $g=g(C)$, with its Jacobian $\mathrm{Jac}(C)$. Mordell-Weil theorem states that $\mathrm{Jac}(C)(K)$ is a finitely generated abelian group, i.e., isomorphic to $\mathbb{Z}^{r}\times G$, where $G$ is a finite group. The number $r$ is called the rank of $\mathrm{Jac}(C)$ over $K$.
@@
Even though there are techniques to compute the rank $r$ in certain cases, for example, using descent [SCH95, STO01, SS03], it is still a difficult problem to determine the exact value of $r$. In principle, the problem can be split into two tasks. The lower bound on $r$ can be obtained from finding enough $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. The upper bound on $r$ can be computed from using descent and computing Selmer groups of $\mathrm{Jac}(C)(K)$. When the two bounds are equal, $r$ can be explicitly determined. However, the bounds are often quite different, making computation of the rank difficult. While there are many families of curves whose ranks of Jacobians are known to be bounded by any side, there are much fewer examples of families of curves with exact ranks.
Hence, for a concrete random curve, the determination of the rank of its Jacobian is a difficult problem. Now, we consider a more challenging question by studying ranks of Jacobians of families of curves, and in our paper, we consider hyperelliptic curves. We adopt the notation from [23]:
[…]
In their recent paper [23], Koymans and Morgan proved a breakthrough result that $1\in\mathcal{R}(K,g)$ for any number field $K$ and any integer $g\geq 1$. The main result of our paper, which we prove in Theorems 2.8 and 3.15 is that $2\in\mathcal{R}(K,g)$ for any number field $K$ and any integer $g\geq 2$.
Theorem 1.1.
Let $g\geq 2$ be an integer and $K$ a number field. There are infinitely many non-isomorphic hyperelliptic curves $C$ of genus $g$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=2$.
Until very recently, it was unknown if $2\in\mathcal{R}(\mathbb{Q},1)$, i.e., if there are infinitely many elliptic curves over $\mathbb{Q}$ of rank exactly 2. This was proven by Zywina in [ZYW25c]. In our paper, as a consequence of our constructions, we obtain a related but strategically different result; namely, in Corollary 2.2, we prove that
$\{0,1,2,3,4,5,6,7,8,9,10,11\}\subseteq\mathcal{R}(\mathbb{Q},2).$
1.1. Motivation
The problem of studying ranks of certain families of elliptic curves has attracted considerable attention in past several years. There are also great mathematical merits of knowing the ranks of Jacobians of curves over number fields, as explained in more detail in §1.4. There are many previous studies which analyse the statistics of rank over infinite family of curves, for example over universal families of elliptic curves [BS15], quadratic twist families of elliptic curves [KMR14, SMI22a, SMI25], cubic twist families of elliptic curves [ABS22, KS24], quartic twist families of elliptic curves over $\mathbb{Q}(i)$ [34], constructions of infinitely many elliptic curves of rank exactly equal to $1$ over any number field [KP25, ZYW25b], and infinitely many elliptic curves of fixed rank between 0 and 4 over any number field [ZYW25a].
More recently, there have been results about the analogous question of ranks of Jacobians of curves of genus $g\geq 2$, but this subarea of research is widely open. While there are some groundbreaking statistical results, such as the results by Yu [YU19] and Smith [SMI22b] who demonstrate infinitely many (respectively positive proportion) of quadratic twist of abelian varieties have rank 0, infinite families of concrete rank remain pretty unknown, especially in the case of a positive rank. The only result we are aware of in this direction is the work by Koymans and Morgan on constructing infinitely many hyperelliptic curves of any genus $g$ [23], defined over any number field $K$, whose Jacobian has rank 1 over $K$. Our work can hence be considered as a continuation of the current progress to construct infinite families of hyperelliptic curves of any genus whose Jacobian has fixed positive rank.
1.2. Our strategy
Our novel and key idea is to intertwine geometry of curves with arithmetic tools to construct new examples of families of hyperelliptic curves of a fixed rank. To elaborate, we produce new hyperelliptic curves by glueing or taking adequate fibre products of two carefully chosen hyperelliptic curves. This strategy is discussed in detail in §3.1, where in the generic case the Jacobian of the new curve is isogenous to a product of Jacobians of three curves of smaller genus. However, as we discuss in §3.2, it is difficult to combine arithmetic tools with the generic case to produce the desired family of hyperelliptic curves. Instead, our choice of curves allow us to bypass this difficulty, by ensuring that the Jacobian of the new curve is isogenous over $K$ to a product of Jacobians of the previously chosen two curves of smaller genus. As far as we know, this geometric strategy has not been used in investigating this question before.
We can then proceed by quadratically twisting the two hyperelliptic curves to approach the problem. Our strategy of the proof comes in two flavours. Either we fix one of the two curves and quadratically twist the other, or we twist both curves with a good control on the rank of both Jacobians concurrently. To ensure the Jacobians of our new curves have fixed positive rank, we combine our geometric strategy with previous results regarding constructions of hyperelliptic curves whose Jacobian has smaller but fixed rank. One subtlety remains in taking this strategy, in that one needs to ensure that the technical conditions appearing in previous results are applicable to our setup. This is rigorously demonstrated in our article, see in particular §3.3. Our article demonstrates that all relevant conditions appearing in such previous works can be checked and satisfied even after undertaking the constructions from the paper for two hyperelliptic curves over $K$.
Notation. Throughout this manuscript, we will not write explicitly the adjective non-isomorphic and the construction of infinitely many curves will always assume they are non-isomorphic. The proof that they are non-isomorphic in each case is straightforward, because we construct the curves by glueing or taking fibre products of twists of hyperelliptic curves, and for infinitely many twists, we introduce a new prime of bad reduction (using which we twist them with square-free elements).
Thin family. We explain why our careful geometric construction for studying ranks of Jacobians of hyperelliptic curves is a subtle non-trivial strategy. Recall that for $g\geq 2$, the dimensions of the moduli space of principally polarised abelian varieties of dimension $g$, the moduli space of curves of genus $g$, and the moduli space of hyperelliptic curves of genus $g$ over $\mathbb{C}$ are respectively given by $\frac{g(g+1)}{2}$, 3g33g-3, and 2g12g-1. For $g\geq 3$, it is clear that the hyperelliptic locus is a thin subset of both moduli spaces of principally polarised abelian varieties and moduli spaces of genus $g$ curves, so the strategy of simply taking products of lower-dimensional abelian varieties does not construct Jacobians of hyperelliptic curves. Even for $g=2$, it is not clear whether a product of two elliptic curves is isogenous to the Jacobian of a hyperelliptic curve over a given number field $K$. These considerations highlight that our technique allows us to obtain rank statistics over a thin subfamily of the moduli space of principally polarised abelian varieties.
1.3. Brief summary of the content of the article
In Section 2, we motivate our general construction of new hyperellitpic curves by first considering bielliptic genus two curves $C$ such that $\mathrm{Jac}(C)$ is isogenous to the product of two elliptic curves $E_{1}$ and $E_{2}$. We divide the construction into 3 cases, depending on the 2-torsion subgroup of $E_{1}$ and $E_{2}$, which is defined over $K$.
  • In §2.1, we consider elliptic curves $E_{1},E_{2}/\mathbb{Q}$ such that $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})=0$, and show that $r\in\mathcal{R}(\mathbb{Q},2)$ for any $0\leq r\leq 11$. In §2.2 and §2.3, in the case $E_{1}[2](\mathbb{Q})\cong E_{2}[2](\mathbb{Q})\neq 0$, we show $r\in\mathcal{R}(\mathbb{Q},2)$ for any $0\leq r\leq 4$ by constructing explicit infinite families of genus 2 curves and rank $r$. These examples are valuable because examples of families with explicit parametrisations are very rare; we only know for a few instances, [FRE84, MON92, DV18].
  • In §2.4, we discuss rational points on these newly constructed genus $2$ hyperelliptic curves.
  • In §2.5, we show that $0,1,2\in\mathcal{R}(K,2)$ over any number field $K$. These constructions use groundbreaking results in [SMI22b, KP25], which prove $0\in\mathcal{R}(K,1)$ [SMI22b] and $1\in\mathcal{R}(K,1)$ [KP25]. We also rigorously demonstrate the applicability of technical assumptions appearing in [KP25] to our setup.
With the genus $2$ case explored, we proceed to Section 3 to consider hyperelliptic curves $C$ of genus $g\geq 3$ such that $\mathrm{Jac}(C)$ is isogenous to the product of two Jacobians of some carefully constructed hyperelliptic curves $C_{1}$ and $C_{2}$.
  • In §3.1, we discuss our geometric strategy of constructing a new hyperelliptic curve by taking fibre products of two hyperelliptic curves. By carefully choosing the two curves, we demonstrate in §3.2 that the Jacobian of the new curve is isogenous to a product of Jacobians of two chosen curves of smaller genus.
  • In §3.3, we show $0,1,2\in\mathcal{R}(K,g)$ for any number field $K$ and any genus $g\geq 3$. The overall strategy of the proof remains identical, but we employ different previous results to generate such curves [YU19, 4, SMI22b, 23]. Technical conditions appearing in [SMI22b] and [23] are once again applicable to our construction of higher genus curves.
  • In §3.4, we further prove $3\in\mathcal{R}(\mathbb{Q},g)$ for $3\leq g\leq 6$ and $4\in\mathcal{R}(\mathbb{Q},g)$ for $3\leq g\leq 4$.
The breakthrough work by Koymans and Morgan [23] in fact answers the previously posed challenge problem, which asks for generating infinitely many absolutely simple Jacobians of hyperelliptic curves over $\mathbb{Q}$ with rank $1$. We mention the formulation of this question in Section 4.
1.4. Importance of Rank(Jac((C)(K)))
@@
  • The rank of $\mathrm{Jac}(C)(K)$ is useful for studying the set $C(K)$, as one can use Chabauty and Coleman method [CHA41, COL85] to determine $C(K)$. Some generalisations of this method can be found in [KIM05, KIM09, BDM+19, SIK09]. These techniques have led to numerous results on studying modular curves that parametrise certain families of elliptic curves, their modularity, torsion subgroups, isogenies and residual Galois representations.
@@
  • The proof of the Hilbert 10th problem over the ring of integers $\mathcal{O}_{K}$, where $K$ is a number field, crucially uses the construction of an abelian variety $A/K$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(A(L))=\mathrm{Rank}(A(K))>0$. This was first proven in the groundbreaking work by Koymans and Pagano [KP24] who constructed an elliptic curve $E/\mathcal{O}_{K}$ and a quadratic extension $L/K$ such that $\mathrm{Rank}(E(L))=\mathrm{Rank}(E(K))=1$. One can also construct infinitely many elliptic curves of rank exactly 1 which are not $\overline{K}$-isomorphic, as proven independently in [ZYW25b]. A different proof of Hilbert 10th problem using positive rank of Jacobians of some twist families of hyperelliptic curves was shown by another groundbreaking work by Alpöge, Bhargava, Ho, and Shnidman [ABH+25].
@@
  • The Birch-Swinnerton-Dyer conjecture [BS65] proposes intricate relations between algebraic properties and analytic properties of abelian varieties. Finding evidence supporting the conjecture has harbored fruitful results in understanding arithmetic properties of elliptic curves [GZ86, GRO91, ZI04, DLR15].
@@
In the case of modular curves, using Eichler-Shimura relations, one may decompose their Jacobians and compute their rank by computing the rank of their components. Still, this technique is not feasible for most curves. Even for hyperelliptic curves, for which we can try to apply the 2-descent, it is not clear if the upper bound obtained from descent will match the number of $\mathbb{Z}$-linearly independent elements in $\mathrm{Jac}(C)(K)$. For example, in [GAJ22], the first named author presents infinite families of curves over $\mathbb{Q}$ whose Jacobians have large rank. Namely if we denote by $g(C)$ the genus of $C$, then an infinite family of hyperelliptic curves such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)-1$ is given in [GAJ22, Theorem 12(1)], whereas $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))\geq g(C)$ is given in [GAJ22, Theorem 12(2)]. The main idea was to find curves with too many rational points that would violate Coleman’s bound [COL85], or Stoll’s improvement [STO06]. However, giving an exact rank of such curves seems extremely difficult.